please solve this
cosx*cosy+sinx*siny=-1 then cosx+cosy

Grade:11

2 Answers

Anurag Singh
6 years ago
cosxcosy+sinxsiny=-1.
=cos(x-y)=-1. >>>>>>>>>>>>equation 1
>>>>  cosx+cosy= 2cos(x+y/2)cos(x-y/2).
but by equation 1 ,x-y=pi(for principle solution).
thus , cos(x-y/2)=cos(pi/2)=0.
so cosx +cosy=2cos(x+y/2)cos(x-y/2)=0
amit dubey
6 years ago
you can get the answer by directing putting value of x and y which satisfy given equation
cosxcosy + sinxsiny = -1
 putting x= 180 and y=0 we get
-1*1 + 0* 0 = -1 
now putting value of x and y in 
cosx+ cos y= -1 + 1= 0